OpenStudy (anonymous):

If the slope of a line is 4/5, does the line rise (from left to right), fall (from left to right), horizontal or vertical?

OpenStudy (anonymous):

can someone plz show me how to solve this, I have three more probs like this

OpenStudy (anonymous):

actually what if the slope is undefined?

OpenStudy (anonymous):

oh wow I really screwd up that assignment

OpenStudy (anonymous):

lol than ks

OpenStudy (anonymous):

wait so i put rises left to right, that's wrong?

sam (.sam.):

no, 

sam (.sam.):

You know the formula of finding a slope is \[m=\frac{y_2-y_1}{x_2-x_1}\] So when m=4/5, \[\frac{4}{5}=\frac{y}{x}\]

OpenStudy (anonymous):

yay so I was right, maybe I didn't screw up!!!!!!!

sam (.sam.):

mhm :)

OpenStudy (amistre64):

might help to visualize a fraction as: the top number is above, or up .... 4 up ----- ; and connect the ends 5 over

OpenStudy (amistre64):

in the case of a negative fraction ... you simply have to adjust the direction you first move :) but as a sue, the top, or up, number moves up and down

OpenStudy (amistre64):

* but as a rule ....